# Spatial Context-based Self-Supervised Learning for Handwritten Text Recognition

### v1

@@ 4.1 Datasets
The proposed methods are assessed using three reference public datasets for HTR: the IAM Handwritten Database [51], the CVL database(CVL) [39], and George Washington (G.W.) [25]. Tab. 1 provides a summary of the characteristics of the corpora at a word-level. For the pre-training, the SSL methods described above use only samples of the training split. We consider the Aachen split [^1] https://www.openslr.org/56/ for the IAM set and that shown in Fischer et al. [25] for Washington. Since CVL does not contain an official validation split, we use 10% of the training data for validation. Note that, in all cases, we transform the input images to grayscale and normalise them to $H=64$, maintaining the aspect ratio.
@@
Images Alphabet Writers Period Language
IAM  [51] 79 275 78 657 Modern Eng
G.W. [25] 4 894 82 2 Historic Eng
CVL  [39] 99 904 53 310 Modern Eng, Ger
The performance of the methods was analysed by employing two widely used text recognition metrics: the Character Error Rate (CER), defined as the normalised Levenshtein distance between ground-truth and prediction, and the Word Accuracy (WAcc).
@@ 4.2 Neural architectures
As mentioned previously, the recognition framework is based on an encoder-decoder scheme. Details of the precise neural architecture used are presented below.
  • Encoder: We considered a slightly modified version of Resnet-34 [30], in which we used fewer channels in intermediate layers and reduced the aggressiveness of the pooling operations, obtaining a similar capacity in terms of parameters as in [16, 6]. We opted not to employ the ResNet-29 model used in the aforementioned works, since this particular architecture was specifically designed for STR tasks—and, more precisely, for addressing images of a fixed size of $32\times 100$—in which the aspect ratio of the images is, unlike that which occurs with HTR, not particularly relevant.
  • Recurrent feature extractor: The sequential nature of the HTR data in question was modelled by considering two stacked bidirectional LSTMs [32] with a hidden size of $T=256$.
  • Transformer decoder: We followed the vanilla decoder implementation used in Vaswani et al. [72], fixing a single decoding layer with 8 attention heads.
With regard to the training details, a maximum number of 1000 epochs with a patience of 50 on the validation loss has been considered for both the pre-training and the downstream stages. Note that, in the latter case, this patience is increased to 200 epochs when a reduced part is used. The weights of the model are optimized using the Adam method [38], fixing the learning rate values to $3\cdot 10^{-4}$ and $10^{-4}$ for the pre-training and fine-tuning phases, respectively. Only the training data is used for the self-supervised pre-training. Data augmentation is also contemplated in both stages, the details of which are provided in the Supplementary material.
@@ 5 Results
The results of our experiments are presented below, and are divided into two main sections. In the first, we analyze the quality of the representations learned for HTR by the different methods. In the second, we compare these methods in a transfer learning scenario, considering cases with both limited and abundant data for fine-tuning.
In all cases, we compare our results with those obtained by ChaCo [81], given that this method has already been shown to represent the state of the art with respect to other proposals described in Section 2.
@@ 5.1 Quality evaluation
The evaluation of SSL methods often hinges on the quality of the representations learned for the task, without the aid of any auxiliary training. This quality is critical, as it directly represents the ability of the model to learn broad and general representations. In order to assess this aspect in the context of HTR, we freeze all representation layers after pre-training so as to then fine-tune them using all the training data. Note that this fine-tuning process is restricted to the final layer that projects the representations onto the character space. By freezing the representation layers, we ensure that the figures reporting recognition accuracy can be attributed to the quality of the representations themselves, rather than enhancements in the architecture of the model or the fine-tuning process as a whole.
Table 5.1 provides details of the quality assessment, averaging across all the datasets considered, [^2] The detailed results for each dataset can be found in the Supplementary material. setting our methods against the ChaCo benchmark. Focusing on our two primary SSL blocks—Geometric and Puzzle—it will be noted that both have some configurations that outperform those of ChaCo, thus demonstrating their utility for HTR tasks. With regard to the best-performing options, the proposed geometric approach “Flip” combined with TD emerges as the most promising alternative, achieving a CER and a WAcc of 17.3 % and 67.8 %, respectively, and clearly outperforming the current state of the art of ChaCo by a noticeable margin (31 % of CER and 53 % of WAcc).
Table 2: Results obtained in the representation quality evaluation in terms of the CER and WAcc figures of merit (%) for both the proposals presented and the reference ChaCo method [81]. The up and down arrows respectively denote whether the metric is positively or negatively valued. The figures represent the average of the individual performances of the datasets contemplated, with the best values per decoding scheme and metric being highlighted in bold type.
CTC TD
CER ($\downarrow$) WAcc ($\uparrow$) CER ($\downarrow$) WAcc ($\uparrow$)
@@
Rotation 27.4 45.3 17.8 67.8
Flip 24.4 52.4 17.3 66.9
@@
Jigsaw 37.3 39.9 35.7 50.3
Sorting 54.7 024.9 57.8 27.6
ChaCo [81] 44.5 33.3 31.1 53.9
@@
Of the results reported, it is worth noting a rather counterintuitive phenomenon: the only method that leverages the sequential nature of the whole word image (“Sorting”) leads to noticeably poorer representations. This outcome is surprising, because one would expect this pretext task to provide a comprehensive view of the input, closely mirroring what the network experiences during the actual HTR task. In contrast, partial images (crops) yield more informative representations for HTR, at least from the point of view of pure pre-training. This reinforces some of the conclusions drawn in literature, which suggest that focusing on local patterns such as single characters or strokes favours the learning process [49, 81]
@@
Concentrating on more general trends, note that the TD setting systematically makes better use of the representations learned through SSL. This trend is evident, as all the TD settings outperform their counterparts that employ CTC. The representations learned would, therefore, appear to reflect an intrinsic compatibility with the TD approach for the decoding of character sequences. Furthermore, the order of methods according to their performance remains consistent regardless of the decoder utilized, whether CTC or TD. This consistency suggests that the quality of the representations is inherently sound across different decoders, a valuable trait for the development of robust HTR systems.
5.2 Transfer learning
We now explore a transfer learning scenario in depth in order to evaluate the adaptability and efficiency of the pre-trained SSL methods. The essence of this experiment is to observe how well these methods can leverage a certain amount of labelled data—5%, 10%, or 100% of the dataset—so as to fine-tune the entire network starting from the representations learned. This fine-tuning is comprehensive, involving all the layers, which contrasts with the previous experiment in which only the final layer was fine-tuned. We aim to discern the ability of these methods to obtain useful representations in HTR as a downstream task and, ultimately, to achieve the highest possible recognition accuracy.
Table 5.2 shows the results of this experiment, providing the average WAcc across all the datasets and the average relative improvements when compared to the baseline which is, in this context, defined as the performance without any pre-training. For the sake of compactness, we provide only the best decoder (CTC or TD) for each method. [^3] The detailed results for each dataset and decoder can be found in the Supplementary material. These metrics provide a holistic view of the performance of each method, thus allowing us to compare not only the absolute accuracy achieved but also the extent to which pre-training contributes to performance gains over the baseline.
Table 3: Average results in terms of the WAcc (%) metric across all datasets for the transfer learning scheme, considering different percentages of labelled data. Both the case in which no pre-training strategy is applied and the reference ChaCo method [81] are included for the purpose of comparison. Values shown in bold type denote the best results obtained per percentage of labelled data, whereas underlined figures highlight the configurations that outperform the ChaCo reference case. The average relative improvement when compared to the No pretrain scenario is presented in parentheses for each case.
Percentage of labeled data
5% 10% 100%
Baseline
No pretrain 15.8 24.1 72.9
ChaCo [81] 24.8 ($\uparrow$ 68.3) 32.9 ($\uparrow$ 46.1) 77.2 ($\uparrow$ 6.5)
[…]
Rotation 24.0 ($\uparrow$ 63.7) 37.7 ($\uparrow$ 78.6) 79.5 ($\uparrow$ 9.4)
Flip 26.0 ($\uparrow$ 75.3) 39.4 ($\uparrow$ 90.0) 79.5 ($\uparrow$ 9.6)
Puzzle 26.5 ($\uparrow$ 103.0) 36.6 ($\uparrow$ 71.9) 78.7 ($\uparrow$ 8.4)
Sorting 15.6 ($\uparrow$ 5.4) 25.3 ($\uparrow$ 12.6) 78.3 ($\uparrow$ 7.9)
In the most data-constrained scenario utilizing merely 5% of the labelled data for fine-tuning, our proposed SSL methods not only clearly surpass the baseline, as one might expect, but also exhibits a slight but consistent advantage over the state-of-the-art ChaCo method. The only exception is “Sorting”, which performs on par with the baseline. With a modest increase to 10 % for fine-tuning, the SSL methods continue to show marked improvements over the baseline, with even greater relative gains. At this level, the proposed methods, save again “Sorting”, begin to distance themselves further from ChaCo in terms of improvement (around 5 % of WAcc).
When all available data is employed for fine-tuning, all competing approaches (including the baseline) surpass 70 % of WAcc. Despite this competitiveness, the SSL methodologies proposed in this work are still those that perform the best, consistently outperforming both the baseline and ChaCo. This new level of performance is also attained by “Sorting”, unlike that which occurred in the previous scenarios, although still below our other proposals.
@@ 5.2 Transfer learning
Although many of our proposals reflect robust behaviour that improves on the state of the art in most cases, the “Flip” method is positioned as the best-performing alternative in general terms. This method not only outperforms others in this semi-supervised scenario but was also superior in the previous section, which focused solely on the quality of the representations learned. Notably, “Flip” is tailored specifically to HTR, thus reinforcing the value of developing and implementing strategies that are directly aligned with the unique challenges and characteristics of the task in hand.
5.3 Consistency
A key objective of our study is to provide practitioners with specific guidelines with which to train more consistent HTR models through the use of SSL. In order to illustrate the effectiveness of the proposed methods, Fig. 4 provides a compelling visual representation that shows the percentage of instances in which any of our methods outperform the state-of-the-art ChaCo for any data scenarios and architectural frameworks (TD or CTC). These bars indicate that, with the exception of the “Sorting” approach, all our methods consistently perform better than that of ChaCo in the majority of cases. The methodologies introduced in this work could, therefore, be adopted as best practices, providing a greater probability of developing robust models for specific scenarios and/or architectures in the context of HTR. The “Flip” method with cropped patches specifically stands out, making it the go-to strategy for those seeking to achieve the most robust performance.
Figure 4: Comparison of occurrences of the proposed methods in the semi-supervised scenario in which they outperform the state-of-the-art ChaCo method [81] for any data collection and architectural framework (TD or CTC). The figures are provided as a ratio with respect to the total number of experimental scenarios considered. The dashed line indicates the chance of our methods to surpass the ChaCo strategy.
@@ 6 Conclusions
In this paper, we study spatial context-based SSL methods originally devised for classification tasks and adapt them to the HTR field. We propose two possible adaptation frameworks for this purpose: (i) adjusting the data in hand in order to use the original SSL methods (input adaptation), and (ii) adapting these SSL approaches to the sequential nature of HTR (method adaptation). The results obtained when considering three benchmark HTR datasets—IAM, CVL and George Washington—show that these proposals not only improve the recognition rates of the model when not considering any pre-training, but also consistently outperform the current state-of-the-art SSL methods for HTR. Moreover, it should be noted that one of our proposals that was specifically designed for HTR (flip) consistently obtains the overall best results in every experimental scenario. Note also that our results are aligned with those of previous SSL literature on text recognition [49], which argues that there is a close relationship between HTR and spatial context owing to stroke and font style information.
The objective of this study is to provide practical guidelines in order to take advantage of SSL techniques in the context of HTR, thus combating the problem of insufficient data in this field.
Appendix A Data augmentation
We detail the data augmentation techniques applied in our study, leveraging the imgaug augmentation package [35]. Our selection is informed by seminal contributions in the field [3, 81], leading to the adoption of the subsequent alterations:

### v2

@@ 4.1 Datasets
The proposed methods are assessed using three reference public datasets for HTR: the IAM Handwritten Database [26], the CVL database (CVL) [21], and RIMES [15]. Table 1 provides a summary of the characteristics of the corpora at a word-level. For the pre-training, the SSL methods described above use only samples of the training split. We consider the Aachen split [^1] https://www.openslr.org/56/ for the IAM set. Since CVL does not contain an official validation split, we use 5% of the training data for validation. In RIMES, we ensure text-letter independence (more details in Supplementary materials). In all cases, we transform the input images to grayscale and normalize them to $H=64$, maintaining the aspect ratio.
@@
Images Alphabet Writers Language
IAM 79 275 78 657 Eng
RIMES 228 358 95 1 300 French
CVL 99 904 53 310 Eng, Ger
For comparative purposes with the related literature [41, 33], performance was assessed in terms of Word Accuracy (WAcc). This figure of merit is defined as the percentage of predicted words that exactly matches the corresponding ground truth.
@@ 4.2 Neural architectures
The details of the considered encoder-decoder recognition scheme are presented below:
  • Encoder: We resorted to the ResNet-29 model used in the works of [9] and [1] to provide feasible comparison, but with slight modifications to adapt it to our input characteristics. More details are provided in the Supplementary material.
  • Recurrent feature extractor: The sequential nature of the HTR data was modelled by considering two stacked bidirectional LSTMs with a hidden size of $T=256$.
  • Transformer decoder: We followed the vanilla decoder used by [35], fixing a single decoding layer with 8 attention heads and a hidden size of 512.
With regard to the training details, a maximum number of 1000 epochs with a patience of 50 on the validation loss has been considered for both the pre-training and the downstream stages. Note that, in the latter case, this patience is increased to 200 epochs when a reduced part of the dataset is used. Only the training data is used for the self-supervised pre-training. Data augmentation is also contemplated in both stages, the details of which are provided in the Supplementary material. The weights of the model are optimized using the Adam method [20], fixing the learning rate values to $3\cdot 10^{-4}$ and $10^{-4}$ for the pre-training and fine-tuning phases, respectively.
@@ 5 Results
The results of our experiments are presented below, and are divided into two main sections. In the first, we analyze the quality of the representations learned for HTR by the different methods. In the second, we compare these methods in a transfer learning scenario, considering cases with both limited and abundant data for fine tuning. In all cases, we compare against state-of-the-art SSL methods for HTR.
@@ 5.1 Quality evaluation
The evaluation of SSL methods often hinges on the quality of the representations learned for the task, without the aid of any auxiliary training. This quality directly represents the ability of the model to learn broad and general representations. To assess this aspect in the context of HTR, we freeze all representation layers after pre-training to then fine-tune the rest of the model using all the training data. By freezing the representation layers, we ensure that the figures reporting recognition accuracy can be attributed to the quality of the representations themselves, rather than enhancements in the architecture of the model or the fine-tuning process as a whole.
Table 2 provides details of the quality assessment for each method and dataset. First, focusing on our two primary SSL blocks—Geometric and Puzzle—it will be noted that the Geometric-based methods show a clear superiority against the Puzzle-based ones while being the last still competitive in several cases, thus demonstrating the utility of spatial information in HTR tasks. Concerning the best-performing options, ChaCo seems to be the most effective in general terms, although the Geometric strategies are close in performance. In this regard, the average differences for each case between the best-performing method among the Baselines and the top Geometric method are 6.2 and 2.8 WAcc points for CTC and autoregressive decoders, respectively.
[…]
Method CTC Autorregresive
IAM CVL RIMES IAM CVL RIMES
Baselines
SeqCLR† 39.7 66.7 63.8 51.9 74.5 79.5
ChaCo† 70.0 75.6 79.4 72.9 77.6 84.9
CMT-Co† 53.1 72.1 66.0 58.2 74.7 74.7
Text-DIAE‡ - - - 71.0 78.1 -
@@
Rotation‡ 65.2 54.9 85.0 67.6 77.2 82.4
Flip‡ 65.9 55.4 82.0 50.1 67.3 57.1
Puzzle
Jigsaw‡ 36.4 18.1 53.9 46.2 46.5 64.7
Sorting‡ 55.0 26.4 66.1 63.3 52.7 84.3
5.2 Transfer learning
We now explore a transfer learning scenario to evaluate the adaptability and efficiency of the pre-trained SSL methods. The essence of this experiment is to observe how well these methods can leverage a certain amount of labelled data—5%, 10%, or 100% of the dataset—to fine tune the entire network starting from the representations learned during pre-training. This fine-tuning process involves all layers, which contrasts with the previous experiment in which only the final layers were updated. We aim to discern the ability of these methods to obtain useful representations in HTR as a downstream task and, ultimately, to achieve the highest possible recognition accuracy. Note that, the same random samples are considered for each method when using a reduced part of a dataset.
Table 3 presents the results of this experiment reporting the WAcc for all datasets and methods. For the spatial context-based SSL strategies, the Geometric approaches, particularly the “Flip" proposal, demonstrate notable superiority. Specifically, this method exhibits a notable advantage over state-of-the-art methods when using a reduced portion of the training data in the IAM and RIMES datasets. For instance, in the IAM dataset with only 10% of the data, “Flip" outperforms the previously best-performing method in each scenario by 11.4 and 11.1 points for the CTC and TD decoders, respectively. Although the CVL dataset does not show this superiority in data-reduced scenarios, “Flip" still surpasses the previous best method for each case when using the full dataset by 7.3 and 4 points for the CTC and TD decoders, respectively. These results highlight the method’s ability to achieve substantial progress in state-of-the-art performance across different scenarios.
On the other hand, it is also worth highlighting the “Rotation" method, which, while slightly trailing “Flip" in performance, still achieves remarkable improvements over the previous state-of-the-art methods in many scenarios. Regarding the Puzzle-based methods, while they remain competitive, they do not demonstrate better performance over either the state-of-the-art or the Geometric-based methods.
@@ 5.2 Transfer learning
Method IAM CVL RIMES
5 10 100 5 10 100 5 10 100
Baselines
SeqCLR 31.2 44.9 76.7 66.0 71.0 77.0 61.8 71.9 90.1
ChaCo 44.3 53.1 79.8 65.6 70.3 78.1 61.3 73.0 89.2
CMT-Co 47.7 56.0 80.1 71.9 74.5 78.2 67.1 75.4 90.1
PerSec-CNN - - 77.9 - - 78.1 - - -
Geometric
Rotation 60.2 66.3 79.0 39.0 52.5 85.3 79.7 84.4 91.4
Flip 61.7 67.4 79.1 40.6 54.6 85.5 79.8 84.2 91.1
Puzzle
Jigsaw 54.5 64.0 77.7 28.7 42.1 83.8 77.6 81.0 90.8
Sorting 56.4 64.7 79.0 31.8 34.7 78.5 74.1 71.5 90.0
(a) CTC decoder.
Method IAM CVL RIMES
5 10 100 5 10 100 5 10 100
Baselines
SeqCLR† 40.3 52.3 79.9 73.1 74.8 77.8 70.9 77.0 92.5
ChaCo† 46.7 55.5 81.4 68.9 74.0 78.2 68.4 77.0 90.6
CMT-Co† 50.4 55.8 81.9 73.6 76.2 78.7 68.8 77.6 91.2
PerSec-CNN† - - 80.8 - - 80.2 - - -
Text-DIAE‡ 49.6 58.7 80.0 47.9 68.5 87.3 - - -
Geometric
Rotation‡ 62.9 68.1 79.6 53.5 66.5 91.2 80.3 84.8 91.7
Flip‡ 62.5 69.8 80.8 52.1 66.3 91.3 79.0 83.9 89.7
Puzzle
Jigsaw‡ 58.5 68.2 80.7 43.5 63.3 90.1 78.9 83.2 92.0
Sorting‡ 59.2 69.2 80.8 43.4 54.2 86.2 80.4 83.9 91.7
(b) Autorregresive decoder.
6 Discussion
In this section, we delve deeper into the analysis and interpretation of the results. For this purpose, we also visualize the feature extraction of each spatial context-based SSL method, compare efficiency of the different neural architectures, and examine the strengths and limitations of the proposed methods.
6.1 Comparison and feature visualization
@@
Considering the results reported, it is worth noting a rather counterintuitive phenomenon: the only method that leverages the sequential nature of the whole word image (“Sorting”) does not lead the best representations. This outcome is surprising, as one would expect this pretext task to provide a comprehensive view of the input, closely mirroring what the network experiences during the actual HTR task. In contrast, partial images (crops) yield more informative representations for HTR, at least from the point of view of pure pre-training. This argument is supported by Fig. 4, which visualizes the feature extraction of the proposed methods as a heat map, highlighting that “Sorting" produces less effective visual features than the Geometric approaches. In addition, note that the “Jigsaw" method, which focuses at sub-character level (just partial views of characters), also shows a clear worse attention to the stroke patterns. This reinforces some of the conclusions drawn in the literature, which suggest that focusing on local patterns such as single characters or strokes favours the learning process [25, 41].
@@
It may be also observed that in our approaches, the TD setting systematically makes better use of the representations learned through SSL. This trend is evident, as all the TD settings outperform their counterparts that employ CTC. The representations learned would, therefore, appear to reflect an intrinsic compatibility with the TD approach for the decoding of character sequences. Furthermore, the order of methods according to their performance remains consistent regardless of the decoder utilized, whether CTC or TD. This consistency suggests that the quality of the representations is inherently sound across different decoders, a valuable trait for the development of robust HTR systems.
(a) Rotation
(b) Flip
(c) Jigsaw
(d) Sorting
Figure 4: Heat map highlighting the regions targeted by the encoder pre-trained with the different spatial context-based SSL methods.
@@
At last, although many of our proposals reflect robust behavior that equal on the state of the art, the “Flip” method is positioned as the best-performing alternative in general terms. This method not only surpasses other spatial context-based techniques but also outperforms prior SSL methods in the semi-supervised scenario. In particular, “Flip" exhibits clear superiority over the “Rotation" strategy when using CTC decoding. We believe that this advantage arises from Flip’s preservation of the horizontal orientation of text, which aligns well with CTC’s horizontal decoding of feature representations. This reinforces the value of developing and implementing strategies that are directly aligned to the unique challenges and characteristics of HTR task.
6.2 Computational architectural requirements
While CRNN architectures with CTC decoding have dominated the state of the art for years, transformer architectures have recently demonstrated exceptional effectiveness in HTR tasks mainly due to their low inductive bias and language model ability. Nevertheless, this advantage comes at the expense of increased computational complexity and resource demands [23]. Given these trade-offs, it is essential to assess the efficiency and practicality of the chosen architecture. Therefore, this section analyzes the computational requirements of the architectures considered in the work and compares them with the state-of-the-art SSL methods.
Table 4: Table showing the input size, GFLOPs, and number of parameters for each method.
Architecture Input size GFLOPs Params
CTC (ours) [64,] 20.4 48.0M
TD (ours) [64,] 20.7 50.7M
[9] CTC [32, 100] 9.7 48.5M
[9] Att. [32, 100] 24.5 49.5M
Text-DIAE [64, 256] 27.6 68.8M
Table 4 presents the Floating Point Operations (FLOPs) and the number of parameters for the architectures used by the different methods for which we were able to replicate the architecture. Focusing first on the architectures employed in this work, as expected, the one incorporating a TD decoder exhibits higher FLOP consumption. However, since the decoder consists of only one layer, the additional computational cost remains minimal—just 0.3 GFLOPs. This makes it a highly favorable trade-off, offering a modest increase in resource usage and parameters while delivering a substantial improvement in performance. In contrast, Text-DIAE, a Transformer-based method, emerges as the most resource-intensive architecture among the evaluated but it does not deliver the best results as shown in Table 3(b), underscoring the fact that simpler architectures and methods can surpass these more complex approaches in performance. This highlights the importance of balancing architectural complexity with efficiency and effectiveness in HTR systems.
On the other hand, when using CTC, our models exhibit more FLOP consumption compared to other CRNN-based methods (note that ChaCo, SeqCLR and CMT-Co share the architecture of [9]) despite having a similar number of parameters. This is because the input image size in our case is, typically, larger. However, this issue is notable mitigated by the inherent computational simplicity of the methods, as we will discuss in next section.
6.3 Strengths and limitations
In recent years, contrastive learning or dual-branch architectures has significantly advanced capabilities in image classification. However, its direct application to HTR is not straightforward. For example, aggressive data augmentation can cause misalignment between positive and negative pairs, leading to suboptimal representations [1]. In contrast, the spatial context-based SSL methods explored in this work avoid such issues. By focusing solely on stroke patterns within the processed images, these methods can leverage far more aggressive data augmentations, such as strong geometric transformations or croppings. This approach enables the network to encounter a wider variety of data, fostering a deeper understanding of their intrinsic characteristics. Additionally, the simplicity of the presented pretext tasks—e.g., classifying an image crop in terms of its rotation angle—reduces computational training costs compared to dual-branch architectures or generative approaches while achieving similar performance rates.
Concerning the limitations, it has also been proven that dual-branch architectures or generative methods show a higher abstraction capacity in the representations by encapsulating more semantic information. In this regard, since the presented proposals exclusively focus on the visual information extracted from the stroke and font style, its applicability to other off-the-shelf tasks—e.g., text segmentation, writer retrieval or image enhancement—remains uncertain. Future works could explore the generalizability of the different SSL methods.
@@ 7 Conclusions
In this paper, we study spatial context-based SSL methods originally devised for classification tasks and adapt them to the HTR field. We propose two possible adaptation frameworks for this purpose: (i) adjusting the data in order to use the original SSL methods (input adaptation), and (ii) adapting these SSL approaches to the sequential nature of HTR (method adaptation). The results obtained when considering three benchmark HTR datasets—IAM, CVL, and RIMES—show that these proposals achieve state-of-the-art performance in various scenarios, particularly in semi-supervised settings. Moreover, among the proposals, it should be noted that one specifically designed for HTR (“Flip”) seems to be specially effective. Also note that our results are aligned with those of previous SSL literature on text recognition [25], which argues that there is a close relationship between HTR and spatial context owing to stroke and font style information. Then, the final objective of this study is to provide practical guidelines to take advantage of a family of SSL techniques in HTR context, thus encouraging the study and application of different SSL frameworks in this field.
Appendix A Experimental set-up
In this section we expand the experimental configuration considered in the work. More precisely, we detail the data augmentation processes, the visual encoder architecture, and the partitioning of the RIMES dataset used.
A.1 Data augmentation
We leverage hard data augmentation techniques enabled by the SSL methods employed, complemented by additional augmentation strategies inspired by seminal works in the field [1, 41]. This approach results in the adoption of the following transformations by using the imgaug augmentation package [19]:
1aug = iaa.Sequential([
2 iaa.SomeOf((1, 8), [
3 # Geometric
4 iaa.ScaleX((0.75, 1.0), mode=’constant’, cval=255),
5 iaa.ScaleY((0.75, 1.25), mode=’constant’, cval=255),
6 iaa.TranslateX(percent=(-0.1, 0.1), mode=’constant’, cval=255),
7 iaa.TranslateY(percent=(-0.15, 0.15), mode=’constant’, cval=255),
8 iaa.Rotate((-5., 5.), mode=’constant’, cval=255),
9 iaa.ShearX((-25, 25), mode=’constant’, cval=255),
10 iaa.PiecewiseAffine(scale=(0.02, 0.04)),
11 # Color
12 iaa.Sharpen(alpha=(0.1, 0.5), lightness=(0.75, 1.25)),
13 iaa.LinearContrast((0.5, 1.5)),
14 # Blur
15 iaa.MotionBlur(k=(5, 7)),
16 iaa.GaussianBlur((1.0, 2.0)),
17 # Noise
18 iaa.AdditiveGaussianNoise(scale=(1, 25))
19 ], random_order=False)
20])
21# Morphological
22num = random.uniform(0, 1)
23if num < 0.33:
24 image = erode(image)
25elif random.uniform(0, 1) < 0.66:
26 image = dilate(image)
27image = aug(image=image)